dy/dx=1+x+y+xy y(-1)=0 then what is the value of funtion y ?

dy/dx=1+x+y+xy
y(-1)=0
then what is the value of funtion y ?

Grade:12

2 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Nikhil

Factorise RHS

RHS=1+x+y(1+x)= (1+x)(1+y)

dy/1+y = (1+x)dx

now integrate both sides

ln(1+y)=x+x2/2 + C

put x=-1 and y=0

0=-1 + 1/2 +C

C=1/2

ln(1+y)=x+x2/2 +1/2

1+y = ex+x2/2 +1/2

y=ex+x2/2 +1/2 -1

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively :Click here to download the toolbar..

vikas askiitian expert
13 years ago

dy/dx = 1+x+y+xy

dy/dx = 1+x +y(1+x)

dy/dx = (1+x)(1+y)

now using variable separation

dy/(1+y) = (1+x)dx

log(1+y) = (x+x2/2)+c

y(-1) = 0 so

c = 1/2

log(1+y) = x+x2/2+1/2

this is the required solution

approve if u like it